# Simple Policy Optimization

### v3

@@ 1 Introduction
Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a) algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments (Berner et al., 2019; Ye et al., 2020). Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded (Wang et al., 2020). The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” (Engstrom et al., 2020). There are also some works that design novel probability ratio clipping methods or objective functions (Queeney et al., 2021; Garg et al., 2021; Markowitz & Staley, 2023; Chen et al., 2018), or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO (Cobbe et al., 2021; Wang et al., 2023).
There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient (Schulman et al., 2017). During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold (Achiam, 2018). Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence (Sun et al., 2022).
However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto\mathbb{R}$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor.
For each trajectory
@@ 3 Method
However, note that if we simply using the above form as the objective for the policy network, this can lead to significant problems. Now Assume that the estimated value of the advantage $\hat{A}(s_{t},a_{t})$ is positive, the policy network would strive to minimize the KL divergence in order to maximize the objective. This tendency would result in the network preferring to make only minor adjustments to the parameters, thereby hindering its ability to learn effectively from historical data.
@@
We will provide intuition about the objective (12), suppose that $d_{\max}=0.01$, and the function curves is shown in Figure 2. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
@@
Considering that the hyperparameter $d_{\max}$ can be set manually and even dynamically changed at any moment during the training process, SPO can benefit significantly from the Corollary 2.2. In summary, the characteristics mentioned above enable SPO to potentially achieve better monotonic policy improvement than PPO, and its performance depends on the value of the hyperparameter $d_{\max}$, which corresponds to the degree of conservativeness in policy iterations.
@@ 4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see Appendix F. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimensionality corresponding to the action space) and the value network (with a 1-dimensional output).
Partial experimental results are shown in Figure 3, with more detailed results provided in the Appendix C. Here, PPO-PENALTY (PPO-P) and PPO-CLIP with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process. In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO.
We also notice that SPO exhibits the highest policy entropy during training, which indicates that SPO has a better exploration capability in the environment. We speculate that this may be a by-product of limiting the KL divergence. In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Subsection 4.4, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
@@ 4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks [^1] Two neural network structures are shown in the appendix F. affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident?
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers (Vaswani et al., 2017) or DM (Diffusion Models) (Ho et al., 2020).
@@ 4.4 Sensitivity Analysis
We conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of SPO, setting it to 0.01, 0.02, and 0.03, respectively. The experimental results are shown in Figure 6.
We can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$, and the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. This result seems to be expected because the KL divergence $d$ satisfies
@@
The cross entropy is non-negative, and it will be a good choice for the agent to reduce the KL divergence by maintaining a higher policy entropy in the training process.
We believe that there is a fundamental difference between SPO and PPO-E, the latter of which introduces an early stopping strategy to forcibly limit the KL divergence, as SPO always exhibits a higher policy entropy than PPO-E when they have a similar level of KL divergence.
@@ 5 Discussion
We propose Simple Policy Optimization (SPO), a first-order algorithm for policy optimization in reinforcement learning. Experimental results in the Atari 2600 environment demonstrate that compared to the traditional PPO algorithm, SPO significantly improves sample efficiency and effectively limits the average KL divergence between the old and current policies, while exhibiting higher policy entropy. SPO strikes a favorable balance between sample complexity, simplicity and wall-time.
In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the average KL divergence between the old and current policies, and this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network.
We have noticed that some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences (Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020; Lee et al., 2023), called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process (Ziegler et al., 2019; Rafailov et al., 2024; Tunstall et al., 2023). It would be intriguing to apply SPO to this process, as the key difference lies in the fact that the old policy is no longer the policy network at the end of the previous training phase, but instead refers to the base reference policy.

### v4

@@ 1 Introduction
Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a) algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments (Berner et al., 2019; Ye et al., 2020). Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded (Wang et al., 2020). The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” (Engstrom et al., 2020). There are also some works that design novel probability ratio clipping methods or objective functions (Queeney et al., 2021; Garg et al., 2021; Markowitz & Staley, 2023; Chen et al., 2018), or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO (Cobbe et al., 2021; Wang et al., 2023). There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient (Schulman et al., 2017). During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold (Achiam, 2018). Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence (Sun et al., 2022). However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto\mathbb{R}$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor. For each trajectory
@@ 4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see Appendix F. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimensionality corresponding to the action space) and the value network (with a 1-dimensional output). Partial experimental results are shown in Figure 3, with more detailed results provided in the Appendix C. Here, PPO-PENALTY (PPO-P) and PPO-CLIP with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process. In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO. We also notice that SPO exhibits the highest policy entropy during training, which indicates that SPO has a better exploration capability in the environment. We speculate that this may be a by-product of limiting the KL divergence. In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Subsection 4.4, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
@@ 4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks [^1] Two neural network structures are shown in the appendix F. affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident? We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers (Vaswani et al., 2017) or DM (Diffusion Models) (Ho et al., 2020).
@@ 4.4 Sensitivity Analysis
We conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of SPO, setting it to 0.01, 0.02, and 0.03, respectively. The experimental results are shown in Figure 6. We can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$, and the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. This result seems to be expected because the KL divergence $d$ satisfies
@@
The cross entropy is non-negative, and it will be a good choice for the agent to reduce the KL divergence by maintaining a higher policy entropy in the training process. We believe that there is a fundamental difference between SPO and PPO-E, the latter of which introduces an early stopping strategy to forcibly limit the KL divergence, as SPO always exhibits a higher policy entropy than PPO-E when they have a similar level of KL divergence.
@@ 5 Discussion
We propose Simple Policy Optimization (SPO), a first-order algorithm for policy optimization in reinforcement learning. Experimental results in the Atari 2600 environment demonstrate that compared to the traditional PPO algorithm, SPO significantly improves sample efficiency and effectively limits the average KL divergence between the old and current policies, while exhibiting higher policy entropy. SPO strikes a favorable balance between sample complexity, simplicity and wall-time. In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the average KL divergence between the old and current policies, and this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network. We have noticed that some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences (Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020; Lee et al., 2023), called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process (Ziegler et al., 2019; Rafailov et al., 2024; Tunstall et al., 2023). It would be intriguing to apply SPO to this process, as the key difference lies in the fact that the old policy is no longer the policy network at the end of the previous training phase, but instead refers to the base reference policy.